# Planning and Learning in Risk-Aware Restless Multi-Arm Bandit Problem

### v1

@@ 2.4 Learning Problem
The parameters of arms in Problem 2.3 may be unknown in various practical applications. Notable examples include a drug discovery problem when a new drug is discovered in a clinical setup Ribba et al., (2020) and a machine maintenance problem for new machines where state transition functions are unknown Ogunfowora and Najjaran, (2023). As a result, the objective in such contexts is to determine a learning policy that converges to the ideal policy (the solution to Problem 2.3) as quickly as possible. Let us assume that the agent interacts with the system for $K$ episodes and let ${\boldsymbol{\pi}}_{k}$ denote a time-dependent learning policy in episode $k$. The performance of a learning policy is measured by Bayesian regret, which quantifies the difference between the policy’s performance and that of an oracle who possesses complete knowledge of the environment and executes an optimal policy $\boldsymbol{\pi}^{\star}$, i.e.,
@@
where the first expectation is calculated from the prior distribution on $\{P^{i}(\cdot|x^{i},a)\}_{i\in\mathcal{N},x^{i}\in\mathcal{X}^{i},a\in\{0,1\}}$, while the second expectation calculated based on the initial states ${{\boldsymbol{x}}_{0}}$ and the learning policy. Bayesian regret is a widely-adopted metric in numerous studies, e.g., Rusmevichientong and Tsitsiklis, (2010); Agrawal and Goyal, (2013); Russo and Van Roy, (2014); Ouyang et al., (2017); Akbarzadeh and Mahajan, (2023). As a result, we define the following problem.
@@ 3.1 Overview of the Risk-neutral RMAB
Whittle, (1988) introduced a priority index policy as a heuristic for Problem 2.3, which has become widely accepted as the conventional method for solving the RMAB Niño-Mora, (2023). This policy is a result of a relaxation of the original optimization problem. Particularly, the (hard) constraint of activating at most $M$ arms at a time is replaced by a relaxation where the expected average number of activated arms per time period is at most $M$, i.e.,
@@
$\max_{{\boldsymbol{\pi}}\in{\boldsymbol{\Pi}}}\mathbb{E}\left[{\boldsymbol{J}}_{{\boldsymbol{x}}_{0}}(\boldsymbol{\pi})\right]-\lambda\left(\mathbb{E}\biggl{[}\dfrac{1}{T}\sum_{t=0}^{T-1}\lVert{\boldsymbol{A}}_{t}\rVert_{1}\bigg{|}{\boldsymbol{X}}_{0}=\boldsymbol{x}_{0}\biggr{]}-M\right)=\sum_{i=1}^{N}\max_{\pi^{i}\in\Pi^{i}_{M}}\bar{J}^{i}_{\lambda,x^{i}_{0}}(\pi^{i})-M\lambda$ (2)
where the policy functions captured in $\Pi_{M}$ are Markovian, i.e. of the form $\pi^{i}:{\mathcal{X}}^{i}\times{\mathcal{T}}\to{\mathcal{A}}$ , and where
$\bar{J}^{i}_{\lambda,x^{i}_{0}}(\pi^{i}):=\mathbb{E}\left[\sum_{t=0}^{T-1}r^{i}\left(X^{i}_{t},A^{i}_{t}\right)-\dfrac{\lambda}{T}\sum_{t=0}^{T-1}A^{i}_{t}\biggm{|}{X^{i}_{0}=x^{i}_{0}}\right].$
For each arm $i$, let $\pi^{i*}_{\lambda}$ denote the optimal policy obtained by dynamic programming on an MDP parametrized by $(\mathcal{X}^{i},\mathcal{A},\allowbreak\{P^{i}(a)\}_{a\in\{0,1\}},\bar{r}^{i}_{\lambda},x_{0}^{i})$, with $\bar{r}^{i}_{\lambda}(x,a):=r^{i}(x,a)-\lambda a/T$. The policies $\pi^{i*}_{\lambda}$ can be assembled to obtain a solution of (2) using $\boldsymbol{\pi}^{*}_{\lambda}:=(\pi_{\lambda}^{1*},\dots,\pi_{\lambda}^{N*})$.
@@
Determining whether a problem is indexable is not immediately apparent, hence researchers have examined various sufficient conditions for indexability (Glazebrook et al.,, 2006; Niño-Mora,, 2007; Akbarzadeh and Mahajan,, 2022). Under certain conditions, the Whittle index policy is optimal (Gittins,, 1979; Weber and Weiss,, 1990; Lott and Teneketzis,, 2000) and in other cases, the Whittle index policy is close to optimal (Glazebrook et al.,, 2006; Niño-Mora,, 2007; Avrachenkov et al.,, 2013; Wang et al.,, 2020; Akbarzadeh and Mahajan,, 2022). General algorithms for computing the Whittle indices are proposed in (Niño-Mora,, 2007; Akbarzadeh and Mahajan,, 2022), while a numerical search is an alternative for computing the indices (Avrachenkov and Borkar,, 2018; Akbarzadeh and Mahajan, 2019a, ).
@@
To establish indexability conditions for Problem 2.3, we first apply the relaxation and decomposition approach described in Section 3.1. Hence, let $\pi^{i*}_{\lambda}$ denote the optimal policy of an arm that maximizes $\bar{D}^{i}_{\lambda,x^{i}_{0}}(\pi^{i})$ which is
@@
We follow the steps presented in Bäuerle and Ott, (2011) to introduce a new augmented arm risk-neutral MDP, which is equivalent to the risk-aware MDP above and can be solved using dynamic programming. Specifically, for each arm $i\in{\cal N}$, we construct a time-dependent MDP $(\bar{\mathcal{X}}^{i},\mathcal{A},\{\bar{{P}}^{i}(a)\}_{a\in\{0,1\}},\{\bar{{r}}_{t}^{i}\}_{t=0}^{T},{\bar{{x}}}^{i}_{0})$ with $\bar{\mathcal{X}}^{i}:=\mathcal{X}^{i}\times\mathcal{S}^{i}$, $\bar{{P}}^{i}(x^{\prime},s^{\prime}|x,s,a):=P^{i}(x^{\prime}|x,a)\mathbb{I}(s^{\prime},s+r^{i}(x,a))$, $\bar{{r}}_{t}^{i}(x,s,a):=\mathbb{I}(t,T-1)U^{i}(s+r^{i}(x,a))-(\lambda/T)a$ and ${\bar{{x}}}_{0}^{i}:=(x_{0}^{i},0)$. The following result can be immediately derived from Theorem 1 of Bäuerle and Rieder, (2014).
@@
Let $\bar{\pi}_{\lambda}^{i*}$ be an optimal Markovian policy for the augmented arm risk-neutral MDP. Then, one can construct an optimal policy for the relaxation of Problem 2.3 using:
$\pi_{\lambda}^{i*}(x_{0:t},a_{0:t-1},t):=\bar{\pi}_{\lambda}^{i*}(x_{t},\sum_{t^{\prime}=0}^{t-1}r^{i}(x_{t^{\prime}},a_{t^{\prime}}),t).$
Namely, $\max_{\pi^{i}\in\Pi_{H}}\bar{D}^{i}_{\lambda,x^{i}_{0}}(\pi^{i})=\bar{D}^{i}_{\lambda,x^{i}_{0}}(\pi_{\lambda}^{i*})$ where $\Pi_{H}$ is the set of all history-dependent policies.
@@ 3.3 A class of indexable arms
Puterman, (2014, Theorem 4.7.4) provides a set of sufficient conditions under which the optimal policy of a risk-neutral MDP is nondecreasing over the state space. The assumptions and the theorem statement for a risk-neutral arm are presented below.
@@
By definition, an arm is indexable if the passive set is nondecreasing in $\lambda$ for every state. We hence use the aforementioned assumption and identify below a class of indexable arms in the risk-aware setting.
@@
For each $i$, the augmented arm risk neutral MDP is indexable if the arm’s original MDP satisfies Assumptions 3, that $q^{i}(x^{\prime}|x,a)$ is nondecreasing in $a$, and that $r^{i}(x,a)=r^{i}(x)$, for all $x,x^{\prime}\in\mathcal{X}^{i}$, and $a\in\mathcal{A}$.
Theorem 5 indicates that if an arm’s MDP satisfies conditions for its optimal risk-neutral policy to be nondecreasing, then its risk-aware policy version is indexable on an augmented state under minor additional assumptions. In Appendix 12, we provide a set of simple structured MDPs that satisfy the assumptions made in Theorem 5.
There are notable practical implications with respect to these assumptions. In the machine maintenance application Glazebrook et al., (2006); Akbarzadeh and Mahajan, 2019b ; Abbou and Makis, (2019), they imply that the state of each machine has a direct impact on its performance and so, a better state results in a higher reward. In addition, the machines deteriorate stochastically if not maintained, with a chance of worsening to any state (generalizing the right skip-free transitions of Glazebrook et al., (2006)). However, the chance of visiting a worse state is higher if the current state is worse, which is often the case in practice. Also, on the other hand, repairs may not always fully fix the issue (generalizing the deterministic transition of Glazebrook et al., (2006) under repairs) but the chance is higher in better states.
The same model also works for patient scheduling in hospitals. These assumptions imply that patients’ conditions can deteriorate stochastically if not attended to, with worse conditions increasing the urgency for intervention. Treatments, while beneficial, may not always fully restore health immediately, reflecting a probabilistic transition towards recovery. This probability is higher if the patient is in a better state. This reflects realistic clinical scenarios where early and continuous care increases the likelihood of reaching an optimal health state Starfield et al., (2005)
@@ 3.4 Computation of Whittle indices
Primarily, there are three approaches to consider. One approach is problem-specific where the Whittle index formula is computed exactly (Jacko and Villar,, 2012; Borkar et al.,, 2018; Yu et al.,, 2018; Wang et al.,, 2019). The other one is the modified adaptive greedy algorithm (Akbarzadeh and Mahajan,, 2022) which works for any indexable RMAB in discounted and average reward setups. And the last one is numerical search which is either through adaptive greedy (Niño-Mora,, 2007), binary search (Qian et al.,, 2016; Akbarzadeh and Mahajan, 2019a, ), or brute-force search.
In this work, we adapt the numerical search approach described in Akbarzadeh and Mahajan, 2019a to compute the Whittle indices using binary search. The procedure of the binary search algorithm is described below and also presented in Appendix 9 in more detail including its computational complexity.
@@ 4 Learning Policy
In this section, we propose a learning algorithm, called Learning Risk-Aware Policy using Thompson Sampling (LRAP-TS), to address Problem 2.4. In this setup, a learning agent aims to minimize the regret in (13) during $K$ episodes of interaction with the system by balancing the exploration-exploitation trade-off (Auer et al.,, 2002).
@@ 4.1 LRAP-TS Algorithm
Algorithm LRAP-TS operates in episodes each with length $T$. It maintains a posterior distribution $\phi^{i}_{k}$ on the dynamics of arm $i$ and keeps track of $N^{i}_{t,k}(x,a)=\sum_{\kappa=1}^{k}\sum_{\tau=0}^{t}\mathds{1}\{(X^{i}_{\kappa,\tau},A^{i}_{\kappa,\tau})=(x,a)\}$ and $N^{i}_{t,k}(x,a,x_{+})=\sum_{\kappa=1}^{k}\sum_{\tau=0}^{t}\mathds{1}\{(X^{i}_{k,\tau},A^{i}_{k,\tau},X^{i}_{k,\tau+1})=(x,a,x_{+})\}.$
@@ 4.2 Regret bound
${\mathcal{R}}(K)\leq 12N^{2}\bar{|\mathcal{X}|}\sqrt{KT(1+\log KT)}.$
@@ 5 Numerical Analysis
We evaluate the risk-aware Whittle index policy through numerical experiments to validate the robustness and efficacy of our models in both planning and learning contexts. Code is available at Anonymous, (2024).
We first describe experiments assessing our model’s performance under various parameter configurations for scheduling problems. All machines (arms) have the same number of states, generate similar rewards, but differ in Markov dynamics.
In the second set of experiments, we evaluate the learning algorithm for LRRMAB, based on advanced breast cancer models created from real-world data in Le, (2016).
Planning: The parameters explored are: time horizon $T\in\{3,4,5\}$, sorted (from worst to best) state spaces $\mathcal{X}$ for each arm from $\{2,3,4,5\}$, number of arms $N\in\{3|{\cal X}|,4|{\cal X}|,5|{\cal X}|\}$, maximum activated arms $\{\lfloor 0.3N\rfloor,\lfloor 0.4N\rfloor,\lfloor 0.5N\rfloor\}$, using the floor operator. In all experiments, all arms are initialized in the best state,
@@
We consider three utility functions for arm $i$, also shown in Fig. 1. For arm $i$, they are
$U^{i}_{\alpha,\tau,o}(J^{i})=\begin{cases}\mathbb{I}(J^{i}-\tau,0)&\text{if }\alpha=1\\ 1-(\tau)^{(1-1/o)}\max(0,\tau-J^{i})^{1/o}&\text{if }\alpha=2\\ (1+e^{-o(1-\tau)})/(1+e^{-o(J^{i}-\tau)})&\text{if }\alpha=3\end{cases}$
where the risk parameters are the target value $\tau\in\{0.5,0.6,0.7\}$ and the order $o\in\{4,8,16\}$.
@@
The first function maximizes the probability of being above $\tau$. The second shows a risk-seeking tendency for total rewards below the threshold. The third combines risk-seeking behavior below the threshold with risk aversion above it.
@@
The per-step reward of each arm is a linear, non-decreasing function of the state, ranging from 00 to $1/T$. State transitions are chosen from structured matrices $\{{\cal P}_{a}(p)\}_{a\in\{0,1\}}$ parameterized by $p\in[0,1/|X|]$ adhering to the assumptions in Theorem  5, reflecting real-world scenarios like machine maintenance or patient scheduling (See Appendix 12). The $p$ values are linearly spaced within $[0.1/|X|,1/|X|]$ across different arms, ensuring scalability with problem complexity.
(a) $(\alpha,\tau)=(1,0.5)$
(b) $(\alpha,\tau,o)=(2,0.5,8)$
(c) $(\alpha,\tau,o)=(3,0.5,8)$
Figure 1: Sample plots of the three utility functions.
@@
Figure 2: Figure (a) shows the distribution of relative improvements in the objective function achieved by our proposed policy compared to the risk-neutral one in $1944$ different setups. Figure (b) illustrates the distribution of total rewards under both risk-aware and risk-free policies for for one of the arms with the setup $T=5$, $|{\cal X}|=2$, $N=10$, $M=3$, $\alpha=1$, $\tau=0.5$. The red line is set at the target $\tau$.
@@
In all experiments, each policy’s performance was evaluated using Monte Carlo simulations, averaging over 100 sample paths. Given the novelty of our approach, we compare the performance of our proposed risk-aware Whittle index policy against the risk-neutral version.
@@
A total of 1944 setups were executed, covering all possible configurations. Fig. 2(a) summarizes the relative improvement in the objective function achieved by our policy compared to the risk-neutral one. On average, our risk-aware policy improved the objective by $11.4\%$ with results ranging from a decrease of $-7.9\%$ to an increase of $54.6\%$. Overall, $82\%$ of trials showed improvement. Fig. 2(b) shows the distribution of total rewards under both policies for a specific setup. The risk-aware policy achieved $97\%$ positive mass compared to $86\%$ for the risk-neutral policy.
Finally, we note the downside of the risk-aware policy—loss in expected total reward. Fig. 2(b) shows the average total reward is 0.62 for the risk-aware policy versus 0.72 for the risk-neutral one. Across the 1944 setups, the risk-aware policy resulted in a drop of $[-11.6\%,3.23\%]$, averaging $-3.1\%$ compared to the risk-neutral policy, which is expected due to the differing optimization objectives.
Learning: This section outlines the experiments conducted for 2.4. We assume all arms follow the same dynamics as described previously, with transition probabilities parameterized by $p^{i}$. The true values of $p^{i}$ are unknown to the learner, who uses LRAP-TS.
Since computing the optimal policy is infeasible for large numbers of arms, even with estimated or true parameters, we use the risk-aware Whittle index policy (Section 3.4) as a proxy. We refer to the algorithm as Risk-Aware Whittle Index Policy using Thompson Sampling (RAWIP-TS).
We consider three models: two from the family $\mathcal{P}$, and one from a real-world healthcare application. For the first two models, we assume $p^{i}\in[0,1/|\mathcal{X}|]$ for arm $i$, and set the initial estimate $\hat{p}^{i}_{0}$ to the median value $0.5/|\mathcal{X}|$. In the third model, each arm represents a patient with Advanced Breast Cancer Le, (2016), where states represent disease progression (DEAD: 0, PROGRESSING: 1, STABLE: 2, RESPONDING: 3). Patients are treated with either Capecitabine or Lapatinib+Capecitabine, with the second treatment being more expensive. Due to limited hospital resources, the problem is modeled as a restless bandit problem with a fixed budget. The rewards for each state are: $r(\text{DEAD})=0$, $r(\text{PROGRESSING})=1/(3T)$, $r(\text{STABLE})=2/(3T)$, $r(\text{RESPONDING})=1/T$.
(a) $(|{\cal X}|,N)=(4,4)$
(b) $(|{\cal X}|,N)=(4,4)$
(c) $(|{\cal X}|,N)=(5,5)$
(d) $(|{\cal X}|,N)=(5,5)$
(e) $(|{\cal X}|,N)=(4,5)$
(f) $(|{\cal X}|,N)=(4,5)$

### v2

@@ 2.4 Learning Problem
The parameters of arms in Problem 2.3 may be unknown in various practical applications. Notable examples include a drug discovery problem when a new drug is discovered in a clinical setup Ribba et al. (2020) and a machine maintenance problem for new machines where state transition functions are unknown Ogunfowora and Najjaran (2023). As a result, the objective in such contexts is to determine a learning policy that converges to the ideal policy (the solution to Problem 2.3) as quickly as possible. Let us assume that the agent interacts with the system for $K$ episodes and let ${\boldsymbol{\pi}}_{k}$ denote a time-dependent learning policy in episode $k$. The performance of a learning policy is measured by Bayesian regret, which quantifies the difference between the policy’s performance and that of an oracle who possesses complete knowledge of the environment and executes an optimal policy $\boldsymbol{\pi}^{\star}$, i.e.,
@@
where the first expectation is calculated from the prior distribution on $\{P^{i}(\cdot|x^{i},a)\}_{i\in\mathcal{N},x^{i}\in\mathcal{X}^{i},a\in\{0,1\}}$, while the second expectation is calculated based on the initial states ${{\boldsymbol{x}}_{0}}$ and the learning policy. Bayesian regret is a widely-adopted metric in numerous studies, e.g., Rusmevichientong and Tsitsiklis (2010); Agrawal and Goyal (2013); Russo and Van Roy (2014); Ouyang et al. (2017); Akbarzadeh and Mahajan (2023). As a result, we define the following problem.
@@ 3.1 Overview of the Risk-neutral RMAB
Whittle (1988) introduced a priority index policy as a heuristic for Problem 2.3, which has become widely accepted as the conventional method for solving the RMAB Niño-Mora (2023). This policy is a result of a relaxation of the original optimization problem. Particularly, the (hard) constraint of activating at most $M$ arms at a time is replaced by a relaxation where the expected average number of activated arms per time period is at most $M$, i.e.,
@@
$\max_{{\boldsymbol{\pi}}\in{\boldsymbol{\Pi}}}\mathbb{E}\left[{\boldsymbol{J}}_{{\boldsymbol{x}}_{0}}(\boldsymbol{\pi})\right]-\lambda\left(\mathbb{E}\biggl{[}\dfrac{1}{T}\sum_{t=0}^{T-1}\lVert{\boldsymbol{A}}_{t}\rVert_{1}\bigg{|}{\boldsymbol{X}}_{0}=\boldsymbol{x}_{0}\biggr{]}-M\right)=\sum_{i=1}^{N}\max_{\bar{\pi}^{i}\in\bar{\Pi}^{i}}\bar{J}^{i}_{\lambda,x^{i}_{0}}(\pi^{i})-M\lambda$ (2)
where the policy functions captured in $\bar{\Pi}^{i}$ are Markovian and of the form $\bar{\pi}^{i}:{\mathcal{X}}^{i}\times{\mathcal{T}}\to{\mathcal{A}}^{i}$, and where
$\bar{J}^{i}_{\lambda,x^{i}_{0}}(\bar{\pi}^{i}):=\mathbb{E}\left[\sum_{t=0}^{T-1}r^{i}\left(X^{i}_{t},A^{i}_{t}\right)-\dfrac{\lambda}{T}\sum_{t=0}^{T-1}A^{i}_{t}\biggm{|}{X^{i}_{0}=x^{i}_{0}}\right].$
For each arm $i$, let $\bar{\pi}^{i*}_{\lambda}$ denote the optimal policy obtained by dynamic programming on an MDP parametrized by $(\mathcal{X}^{i},\mathcal{A},\allowbreak\{P^{i}(a)\}_{a\in\{0,1\}},\bar{r}^{i}_{\lambda},x_{0}^{i})$, with $\bar{r}^{i}_{\lambda}(x,a):=r^{i}(x,a)-\lambda a/T$. The policies $\pi^{i*}_{\lambda}$ can be assembled to obtain a solution of (2) using $\bar{\boldsymbol{\pi}}^{*}_{\lambda}:=(\bar{\pi}_{\lambda}^{1*},\dots,\bar{\pi}_{\lambda}^{N*})$.
@@
Determining whether a problem is indexable is not immediately apparent, hence researchers have examined various sufficient conditions for indexability (Glazebrook et al., 2006; Niño-Mora, 2007; Akbarzadeh and Mahajan, 2022). Under certain conditions, the Whittle index policy is optimal (Gittins, 1979; Weber and Weiss, 1990; Lott and Teneketzis, 2000) and in other cases, the Whittle index policy is close to optimal (Glazebrook et al., 2006; Niño-Mora, 2007; Avrachenkov et al., 2013; Wang et al., 2020; Akbarzadeh and Mahajan, 2022). General algorithms for computing the Whittle indices are proposed in (Niño-Mora, 2007; Akbarzadeh and Mahajan, 2022), while a numerical search is an alternative for computing the indices (Avrachenkov and Borkar, 2018; Akbarzadeh and Mahajan, 2019a).
@@
To establish indexability conditions for Problem 2.3, we first apply the relaxation and decomposition approach described in Section 3.1. Hence, let $\pi^{i*}_{\lambda}$ denote the optimal policy of an arm that maximizes $\bar{D}^{i}_{\lambda,x^{i}_{0}}(\bar{\pi}^{i})$ which is
@@
We follow the steps presented in Bäuerle and Ott (2011) to introduce a new augmented arm risk-neutral MDP, which is equivalent to the risk-aware MDP above and can be solved using dynamic programming. Specifically, for each arm $i\in{\cal N}$, we construct a time-dependent MDP $(\bar{\mathcal{X}}^{i},\mathcal{A},\{\bar{{P}}^{i}(a)\}_{a\in\{0,1\}},\{\bar{{r}}_{t}^{i}\}_{t=0}^{T},{\bar{{x}}}^{i}_{0})$ with $\bar{\mathcal{X}}^{i}:=\mathcal{X}^{i}\times\mathcal{S}^{i}$, $\bar{{P}}^{i}(x^{\prime},s^{\prime}|x,s,a):=P^{i}(x^{\prime}|x,a)\mathbb{I}(s^{\prime},s+r^{i}(x,a))$, $\bar{{r}}_{t}^{i}(x,s,a):=\mathbb{I}(t,T-1)U^{i}(s+r^{i}(x,a))-(\lambda/T)a$ and ${\bar{{x}}}_{0}^{i}:=(x_{0}^{i},0)$. The following result can be immediately derived from Theorem 1 of Bäuerle and Rieder (2014).
@@
Let $\tilde{\pi}_{\lambda}^{i*}$ be an optimal Markovian policy for the augmented arm risk-neutral MDP. Then, one can construct an optimal policy for the relaxation of Problem 2.3 using:
$\bar{\pi}_{\lambda}^{i*}(x_{0:t},a_{0:t-1},t):=\tilde{\pi}_{\lambda}^{i*}(x_{t},\sum_{t^{\prime}=0}^{t-1}r^{i}(x_{t^{\prime}},a_{t^{\prime}}),t).$
Namely, $\max_{\bar{\pi}^{i}\in\bar{\Pi}_{H}}\bar{D}^{i}_{\lambda,x^{i}_{0}}(\bar{\pi}^{i})=\bar{D}^{i}_{\lambda,x^{i}_{0}}(\bar{\pi}_{\lambda}^{i*})$ where $\bar{\Pi}_{H}$ is the set of all history-dependent policies.
@@ 3.3 A class of indexable arms
Puterman (2014, Theorem 4.7.4) provides a set of sufficient conditions under which the optimal policy of a risk-neutral MDP is nondecreasing over the state space. The assumptions and the theorem statement for a risk-neutral arm are presented below.
@@
For each $i$, the augmented arm risk neutral MDP is indexable if the arm’s original MDP satisfies Assumptions 3, that $q^{i}(x^{\prime}|x,a)$ is nondecreasing in $a$, that $r^{i}(x,a)=r^{i}(x)$, for all $x,x^{\prime}\in\mathcal{X}^{i}$, and $a\in\mathcal{A}$, and that the utility function $U^{i}$ is a nondecreasing function.
Proof Sketch. By definition, an arm is indexable if the passive set is nondecreasing in $\lambda$ for every state. Thus, it suffices to show that the optimal policy of an arm is nondecreasing in $\lambda$. Lemma 4 provides a set of sufficient conditions under which the optimal policy is nondecreasing in $\mathcal{X}$. We adapt these conditions to ensure that this property extends over all possible values of $\lambda$.
Theorem 5 indicates that if an arm’s MDP satisfies conditions for its optimal risk-neutral policy to be nondecreasing, then its risk-aware policy version is indexable on an augmented state under minor additional assumptions. In Appendix F, we provide a set of simple structured MDPs that satisfy the assumptions made in Theorem 5.
There are notable practical implications with respect to these assumptions. In the machine maintenance application Glazebrook et al. (2006); Akbarzadeh and Mahajan (2019b); Abbou and Makis (2019), they imply that the state of each machine has a direct impact on its performance and so, a better state results in a higher reward. In addition, the machines deteriorate stochastically if not maintained, with a chance of worsening to any state (generalizing the right skip-free transitions of Glazebrook et al. (2006)). However, the chance of visiting a worse state is higher if the current state is worse, which is often the case in practice. Also, on the other hand, repairs may not always fully fix the issue (generalizing the deterministic transition of Glazebrook et al. (2006) under repairs) but the chance is higher in better states.
The same model also works for patient scheduling in hospitals. These assumptions imply that patients’ conditions can deteriorate stochastically if not attended to, with worse conditions increasing the urgency for intervention. Treatments, while beneficial, may not always fully restore health immediately, reflecting a probabilistic transition towards recovery. This probability is higher if the patient is in a better state. This reflects realistic clinical scenarios where early and continuous care increases the likelihood of reaching an optimal health state Starfield et al. (2005)
@@ 3.4 Computation of Whittle indices
Primarily, there are three approaches to consider. One approach is problem-specific where the Whittle index formula is computed exactly (Jacko and Villar, 2012; Borkar et al., 2018; Yu et al., 2018; Wang et al., 2019). The other one is the modified adaptive greedy algorithm (Akbarzadeh and Mahajan, 2022) which works for any indexable RMAB in discounted and average reward setups. And the last one is numerical search which is either through adaptive greedy (Niño-Mora, 2007), binary search (Qian et al., 2016; Akbarzadeh and Mahajan, 2019a), or brute-force search.
In this work, we adapt the numerical search approach described in Akbarzadeh and Mahajan (2019a) to compute the Whittle indices using binary search. The procedure of the binary search algorithm is described below and also presented in Appendix C in more detail including its computational complexity.
@@ 4 LEARNING POLICY
In this section, we propose a learning algorithm, called Learning Risk-Aware Policy using Thompson Sampling (LRAP-TS), to address Problem 2.4. In this setup, a learning agent aims to minimize the regret in (13) during $K$ episodes of interaction with the system by balancing the exploration-exploitation trade-off (Auer et al., 2002).
@@ 4.1 LRAP-TS Algorithm
Algorithm LRAP-TS operates in episodes each with length $T$. It maintains a posterior distribution $\phi^{i}_{k}$ on the dynamics of arm $i$ and keeps track of $N^{i}_{k,t}(x,a)=\sum_{\kappa=1}^{k}\sum_{\tau=0}^{t}\mathds{1}\{(X^{i}_{\kappa,\tau},A^{i}_{\kappa,\tau})=(x,a)\}$ and $N^{i}_{k,t}(x,a,x_{+})=\sum_{\kappa=1}^{k}\sum_{\tau=0}^{t}\mathds{1}\{(X^{i}_{k,t},A^{i}_{k,t},X^{i}_{k,t+1})=(x,a,x_{+})\}.$
@@
(a) $(\alpha,\tau)=(1,0.5)$
(b) $(\alpha,\tau,o)=(2,0.5,8)$
(c) $(\alpha,\tau,o)=(3,0.5,8)$
Figure 1: Sample plots of the three utility functions.
@@ 4.2 Regret bound
${\mathcal{R}}(K)\leq 12N^{2}TR_{\max}\bar{|\mathcal{X}|}\sqrt{KT(1+\log KT)}.$
@@ 5 NUMERICAL ANALYSIS
We evaluate the risk-aware Whittle index policy through numerical experiments to validate the robustness and efficacy of our models in both planning and learning contexts. We consider two applications, machine maintenance, and patient scheduling.
We first describe planning experiments assessing our model’s performance under various parameter configurations for a machine maintenance problem instance which is generated synthetically and aligns with Akbarzadeh and Mahajan (2019b). In the second set of experiments, we evaluate the learning algorithm for LRRMAB, based on both the synthetic machine maintenance model and patient scheduling for the advanced breast cancer models created from real-world data in Le (2016).
@@
Within these experiments, we consider three utility functions for arm $i$, also shown in Fig. 1. For arm $i$, they are
$U^{i}_{\alpha,\tau,o}(J^{i})=\begin{cases}\mathbb{I}(J^{i}-\tau,0)&\text{if }\alpha=1\\ 1-\tau^{-1/o}\max(0,\tau-J^{i})^{1/o}&\text{if }\alpha=2\\ (1+e^{-o(1-\tau)})/(1+e^{-o(J^{i}-\tau)})&\text{if }\alpha=3\end{cases}$
@@
where the risk parameters are the target value $\tau$, and the order $o$. The first function maximizes the probability of being above $\tau$. The second shows a risk-seeking tendency for total rewards below the threshold. The third combines risk-seeking behavior below the threshold with risk aversion above it.
The code for all experiments is available on GitHub. [^1] https://github.com/nima-akbarzadeh/AISTATS25
Planning: We consider a machine maintenance scheduling problem with a fleet of non-homogeneous machines (arms) where all of them have the same number of states, generate similar rewards, but differ in Markov dynamics.
@@
The per-step reward of each arm is a linear, non-decreasing function of the state, ranging from 00 to $1/T$. This assumption makes the total reward of a round be bounded by $1$. State transitions are chosen from structured matrices $\{{\cal P}_{a}(p)\}_{a\in\{0,1\}}$ parameterized by $p$ which are linearly spaced within $[0.1/|X|,1/|X|]$ across different arms adhering to the assumptions in Theorem  5 and reflecting the type of dynamics encountered in a machine maintenance problem (See Model $4$ in Appendix F).
[…]
@@
Figure 2: Figure (a) shows the distribution of relative improvements in the objective function achieved by our proposed policy compared to the risk-neutral one in $6804$ different setups. This histogram is limited to range of values up to $100$. Figure (b) illustrates the distribution of total rewards under both risk-aware and risk-free policies for for one of the arms with the setup $T=5$, $|{\cal X}|=5$, $N=25$, $M=7$, $\alpha=1$, $\tau=0.5$. The red line is set at the target $\tau$.
@@
In all experiments, each policy’s performance was evaluated using Monte Carlo simulations, averaging over 100 sample paths. Given the novelty of our approach, we compare the performance of our proposed risk-aware Whittle index policy against the risk-neutral version and provide a relative improvement of the proposed method with respect to the risk-neutral version measured in percentages.
@@
Fig. 2(a) summarizes the relative improvement in the objective function achieved by our policy in a histogram. Some statistics are reported in Table 1. Fig. 2(b) shows the distribution of total rewards under both policies for a specific setup. The risk-aware policy achieved $100\%$ positive mass compared to $58\%$ for the risk-neutral policy.
Mate et al. (2021) addressed a binary state partially-observable restless bandit problem, while our work considers a fully-observable setup. Although the two setups differ fundamentally, we include a numerical baseline inspired by Mate et al.’s who maximize the expected sum of stage-wise utilities instead of the expected utility of the sum of stage-wise performances. To this end, we will include a baseline (named Sum of Stage-wise Utilities Policy (SSUP)) where the objective is modified to: $\sum_{t=0}^{T-1}\frac{1}{T}U^{i}_{\alpha,\tau,o}(Tr^{i}_{t})$. Similar to RAWIP, we compute the relative improvement of this baseline with respect to the risk-neutral Whittle index policy. Results (Table 1) indicate that our risk-aware Whittle index policy outperforms the baseline in all the provided statistics, highlighting its superior performance.
Policy Min (%) Max (%) Mean (%) % Above 0
RAWIP -40.82 4842.0 37.9 84.9
SSUP -45.84 560.1 3.46 64.1
Table 1: Relative improvement of policies compared to the risk-neutral Whittle index policy.
To provide better clarity, we plan to include tables illustrating performance variations with respect to both the threshold ($\tau$) and the utility functions considered in the study. This will help highlight how different utility functions affect performance under various parameter configurations. Please refer to the table below.
Parameter Value Relative Improvement
$\alpha=1$ 120.39%
$(\alpha,o)=(2,4)$ 6.66%
$(\alpha,o)=(2,8)$ 16.89%
$(\alpha,o)=(2,16)$ 34.04%
$(\alpha,o)=(3,4)$ 3.76%
$(\alpha,o)=(3,8)$ 23.73%
$(\alpha,o)=(3,16)$ 59.61%
Table 2: The effect of the utility function on the relative improvement of RAWIP with respect to the risk-neutral Whittle index policy
(a) $(|{\cal X}|,N)=(3,4)$
(b) $(|{\cal X}|,N)=(3,4)$
(c) $(|{\cal X}|,N)=(5,4)$
(d) $(|{\cal X}|,N)=(5,4)$
(e) $(|{\cal X}|,N)=(5,5)$
(f) $(|{\cal X}|,N)=(5,5)$